OpenStudy (anonymous):

What equivalent to 9/20 ?? Answers choices •15/20 •45% •4:10 •12 to 32

OpenStudy (goldphenoix):

Do you think 9/20 is equal to 15/20?

OpenStudy (anonymous):

I really am not sure

OpenStudy (goldphenoix):

Let me restate that. So let say my friend and me were having an eating contest. Whoever can eat 20 pizzas win. I only ate 9 pizzas, but my friend ate 15 pizzas. Does that mean that my friend and me ate the same amount of pizza?

OpenStudy (anonymous):

Do you know the Anwser ?

OpenStudy (anonymous):

\[\frac{9}{20} = \frac{9\times5}{20\times5}= \frac{45}{100}\] = 45%